If the market rate of interest is 10%, a $15000, 13%, 10-year bond that pays interest annually would sell at an amount

equal to face value.

less than face value.

that cannot be determined.

greater than face value.
Answer #1

The correct answer is :- Greater than face value.

------------------------------------------------------------------------------------------------------

Explanation : - When a bonds yield to maturity is less than the coupon rate, the bond sells at a price greater than the face value.

The price of the bond is found using the following equation

Bond price = 2 (1+r):+ (1+r)- ŚC+M

C = Annual coupon payment = 0.13 \times $ 15,000 =$ 1950

M = face value = $ 15,000

r = market rate of interest = 10%

$1950 $15,000 Bond price = 201 +0.10 10 (1 + 0.10)10

Price of the bond = $ 17,765.10